Estimate the rate of heat loss from a vertical thin plate immersed in CO2 at one atmosphere and 10℃. The plate is 4 m high and 3 m wide. It is maintained at 44℃.
Answer #1

Question: biveni Com 11 Ambient temperature ature To=100 Toos 10 + 273 Tx= 283K 3 Plate temperature Ts=446 Ts= 44+273 Ts= 317al 270 for Coo » Density 8 = 1.78 kg/m3 Dynamic 1.5x10-5 Pars > Thermal Conductivity ke 0.0165 WIMK & Prandte number Pr 0.76COM २१५673 ११ ५१8.88 --- Roy leigh number (Raj Ra= Gg x Pr Ra= २१467397498 880.76 Ra= २२ 395 २२२ ०११, 148 Ra= २२ , 395 ₹०१Ra 7 9x109 - Turbulent flow Nusselt number (NO) 다. Nu 0 13 x ( Ra 73 in Nu= 366.433 7 ha Convective heat transfer Coefficientb Nu XK C L h = 366.434 0:0165 D 4 ㅋ h= 1.5115 шI m 2k b= 1:51 W/m2K > Heat Loss (o БА AT =A= Area of plate Q = 151x (4x34 ( 44-10) 여 TT 8 - 61 6,08 Watt - 0.616 Kw